# Asymptotics of the maximum likelihood estimator of the location parameter of Pearson Type VII distribution

### v1

@@ 1. Introduction
By the Lebesgue convergence theorem, $G_{m}^{\prime}$ is continuous. Hence, it suffices to show that $G_{m}^{\prime}(0)<0$.
@@
We also deal with the derivatives of $D_{n}(t)$ and $G_{m}(t)$ with respect to $t$. The following corresponds to [2, (3.32)].
Lemma 3.3.
@@
$\lim_{n\to\infty}\max_{t\in[-r_{m,2},r_{m,2}]}\left|D_{n}^{\prime}(t)-G_{m}^{\prime}(t)\right|=0,\ \textup{ $P$-a.s.}$
@@
By (3.7) below, $|\partial_{t}^{2}D(x,t)|\leq 3$, and hence, the map $t\mapsto\partial_{t}D(x,t)$ is Lipschitz continuous with the Lipschitz constant $3$. Let $Y^{\prime}_{i}:=\partial_{t}D(X_{i},t)-G_{m}^{\prime}(t)$. Since $|\partial_{t}D(X_{i},t)|\leq 1$ and $|G_{m}^{\prime}(t)|\leq 1$, $\left(Y^{\prime}_{i}\right)_{i}$ are i.i.d., $|Y_{i}^{\prime}|\leq 2$ and $E[Y_{i}^{\prime}]=0$. Therefore, we can show this assertion as in the proof of Lemma 3.1. ∎
@@
$P$-a.s. $\omega$, there exists $N_{5}(\omega)\in\mathbb{N}$ such that for every $n>N_{5}(\omega)$, $|C_{Q_{n}(\omega)}\cap[-r_{m,2},r_{m,2}]|=1$.
@@
in particular, $D_{n}(t)(\omega)$ is strictly decreasing in $t$ on $r_{m,2},r_{m,2}]$.
@@
$\mathcal{A}_{n}:=\mathcal{A}_{n,1}\cap\mathcal{A}_{n,2}\cap\mathcal{A}_{n,3}.$
Let $\widetilde{\mathcal{A}_{i}}:=\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n,i},i=1,2,3$.
By Lemma 3.1, $P\left(\widetilde{\mathcal{A}_{1}}\right)=1$. By Lemma 3.3, $P\left(\widetilde{\mathcal{A}_{2}}\right)=1$. By Propositions 2.7 and 3.4, $P\left(\widetilde{\mathcal{A}_{3}}\right)=1$. Since $\widetilde{\mathcal{A}_{1}}\cap\widetilde{\mathcal{A}_{2}}\cap\widetilde{\mathcal{A}_{3}}=\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}$, $P\left(\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}\right)=1$, and in particular, $\lim_{n\to\infty}P(\mathcal{A}_{n})=1$.
@@
$\lim_{n\to\infty}P\left(D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right)=\int_{-\infty}^{y}\varphi_{m}(t)dt,$ (3.1)
where $\varphi_{m}$ is the density function of the distribution $N\left(0,\frac{m+1}{m(2m-1)}\right)$.
@@
By the symmetry,
$E[D(X_{1},0)]=c_{m}\int_{\mathbb{R}}\frac{x}{(1+x^{2})^{1+m}}dx=0.$ (3.2)
@@
$E\left[D(X_{1},0)^{2}\right]=c_{m}\int_{\mathbb{R}}\frac{x^{2}}{(1+x^{2})^{2+m}}dx=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}=\frac{2m-1}{4m(m+1)},$ (3.3)
@@
$\sqrt{n}D_{n}\left(0\right)\Rightarrow N\left(0,\frac{2m-1}{4m(m+1)}\right),\ n\to\infty,$ (3.4)
@@
$E\left[\partial_{t}D(X_{1},0)\right]$ $=c_{m}\int_{\mathbb{R}}\frac{x^{2}-1}{(1+x^{2})^{2+m}}dx$
$=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}-\frac{B(1/2,m+3/2)}{B(1/2,m-1/2)}=-\frac{2m-1}{2(m+1)}.$ (3.5)
@@
$\lim_{n\to\infty}\frac{y}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0)=-\frac{2m-1}{2(m+1)}y,\ \textup{ $P$-a.s.}$ (3.6)
Since
$\partial_{t}^{2}D(x,t)=\frac{2(x-t)((x-t)^{2}-3)}{(1+(x-t)^{2})^{3}},$ (3.7)
$\max_{x,t\in\mathbb{R}}\left|\partial_{t}^{2}D(x,t)\right|=\max_{y\in\mathbb{R}}\frac{2|y||y^{2}-3|}{(1+y^{2})^{3}}=:C_{1}<\infty.$
@@
$\left|D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right|\leq C_{1}\frac{|y|^{2}}{n}.$ (3.8)
@@
$\lim_{n\to\infty}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right)=0,\ \textup{ $P$-a.s.}$ (3.9)
By (3.4), (3.6), (3.9) and Slutsky’s theorem,
@@
We see that (3.1) holds.
@@
We follow the strategy of [3, Section 5]. As in the case of Theorem 1.2, we cannot apply the result of [3] directly and need to modify several parts.
Let $\phi(n):=\sqrt{2n\log\log n}$. Let
$R_{n}:=\frac{1}{\phi(n)}\sum_{i=1}^{n}D(X_{i},0),$
$S_{n}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0),$
@@
$T_{n}^{(\gamma)}:=\frac{1}{\phi(n)}\sum_{i=1}^{n}\left(D\left(X_{i},\gamma\frac{\phi(n)}{n}\right)-D(x_{i},0)-\gamma\frac{\phi(n)}{n}\partial_{t}D(X_{i},0)\right).$
Then,
$R_{n}+\gamma S_{n}+T_{n}^{(\gamma)}=\frac{1}{\phi(n)}\sum_{i=1}^{n}D\left(X_{i},\gamma\frac{\phi(n)}{n}\right).$ (4.1)
Since $\max_{x,t\in\mathbb{R}}|D(x,t)|\leq 1/2$, by the Kolmogorov law of the iterated logarithm, there exists an event $\Omega_{4}$ such that $P(\Omega_{4})=1$ and for every $\omega\in\Omega_{4}$,
$\limsup_{n\to\infty}R_{n}(\omega)=\sqrt{\frac{2m-1}{4m(m+1)}}.$ (4.2)
@@
$\lim_{n\to\infty}S_{n}(\omega)=-\frac{2m-1}{2(m+1)}.$ (4.3)
Let $\Omega_{6}:=\Omega_{4}\cap\Omega_{5}\cap\left(\cup_{N\geq 1}\cap_{n\geq N}\mathcal{A}_{n}\right)$. Then, $P\left(\Omega_{6}\right)=1$.
@@
$\lim_{n\to\infty}T_{n}^{(\gamma)}(\omega)=0.$ (4.4)
For notational convenience, let $\sigma_{m}:=\sqrt{\dfrac{m+1}{m(2m-1)}}$. Let $\omega\in\Omega_{6}$ and $\epsilon>0$. Then, there exists $N_{9}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{9}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)<(\sigma_{m}+\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}+\epsilon)\frac{\phi(n)}{n}\right)<0$ holds. By (4.1), this is equivalent to
$R_{n}(\omega)+(\sigma_{m}+\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}+\epsilon)}(\omega)<0.$ (4.5)
@@
$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\leq\sigma_{m}.$ (4.6)
We can show the lower bound in the same manner. There exists $N_{11}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{11}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)>(\sigma_{m}-\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}-\epsilon)\frac{\phi(n)}{n}\right)>0$ holds. By (4.1), this is equivalent to
$R_{n}(\omega)+(\sigma_{m}-\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}-\epsilon)}(\omega)>0.$ (4.7)
@@
$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\geq\sigma_{m}.$ (4.8)
@@
This completes the proof.
@@
We prove Theorem 1.4 by following the strategy of [2]. In the above section, we have seen that for $i=1,2$, $P(\mathcal{A}_{n,i}^{c})$ decays exponentially fast. Here we show that $P\left(\mathcal{A}_{n,3}^{c}\right)$ also decays exponentially fast.
Recall the definition of $F_{m}$ in (2.1). For notational convenience, let $\widetilde{F}_{m}(t):=\exp\left(F_{m}(t)\right)$.
@@
The following corresponds to [2, (3.15)].
@@
Let $r>0$. Assume that
$0<\delta<\min\left\{\frac{1}{2}\left(m-\frac{1}{2}\right),\frac{F_{m}(r)-F_{m}(0)}{4}\right\}.$
Then, there exists a positive constant $c_{m,3}$ depending only on $m$ such that for every $t$ with $|t|>r$ and every $n\geq 1$,
@@
We assume that $t>r$. The proof is the same for the case that t<rt<-r. We see that
@@
Assume that $0<\lambda<m-1/2$. Then,
$E\left[\exp\left(\lambda\left|F_{m}(t)-\log(1+(X_{1}-t)^{2})\right|\right)\right]\leq\exp(\lambda F_{m}(t))E\left[(1+(X_{1}-t)^{2})^{\lambda}\right]<\infty.$
@@
$E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]=\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right].$
Since $E\left[\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right]=0$,
$\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]=1+\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right].$
@@
$E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq c_{m,4}^{k}P(X_{1}\leq t/2)+F_{m}(t)^{k}P(X_{1}>t/2).$
@@
$E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq c_{m,4}^{k}+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{k},$
where we let $c_{m,5}:=c_{m}4^{m}$. Hence,
$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]$
$\leq\frac{\lambda^{2}}{2}\left(c_{m,4}^{2}\exp(\lambda c_{m,4})+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda F_{m}(t))\right).$
Since $0<\lambda<m-\frac{1}{2}$,
@@
Let $\lambda_{m}:=\frac{1}{2}(m-\frac{1}{2})$. Then, for every $\lambda\in(0,\lambda_{m})$,
$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq\lambda^{2}c_{m,6},$
@@
$c_{m,6}:=\frac{1}{2}\left(c_{m,4}^{2}\exp(\lambda_{m}c_{m,4})+\sup_{t\geq 0}\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda_{m}F_{m}(t))\right)<\infty.$
@@
Therefore, for every $\lambda\in\left(0,\min\left\{\lambda_{m},\frac{F_{m}(r)-F_{m}(0)}{4}\right\}\right)$,
@@
If we let $\lambda:=\delta/c_{m,6}$, then, $0<\lambda<m-1/2$, and,
@@
Let $\lambda_{m}(r):=\frac{1}{2}\min\left\{\frac{1}{2}\left(m-\frac{1}{2}\right),\frac{F_{m}(r)-F_{m}(0)}{4}\right\}$.
The following corresponds to [2, (3.21)] [^4] There is a typo in [2, (3.21)]. The supremum in [2, (3.21)] should be the infimum..
@@
Let $r>0$. Assume that $0<\delta<\lambda_{m}(r)$. Then, there exists $N(r,\delta)\in\mathbb{N}$ such that for every $n\geq N(r,\delta)$,
@@
We remark that $r>0$ can be taken arbitrarily small.
@@
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)\leq\exp\left(-\frac{\delta^{2}}{8c_{m,3}}n\right).$ (5.2)
Since $L_{n}^{\prime}(t)=-2D_{n}(t)$ and $|D_{n}(t)|\leq 1/2$,
@@
The following corresponds to [2, (3.25)] [^5] There is also a typo in [2, (3.25)]. $n^{2}$” in the right hand side of the inequality in [2, (3.25)] should be $n\delta^{2}$”.. Recall that $\lambda_{m}=\frac{1}{2}(m-\frac{1}{2})$.

### v2

@@ 1. Introduction
By the Lebesgue convergence theorem again, we see that $G_{m}^{\prime}$ is continuous. Hence, it suffices to show that $G_{m}^{\prime}(0)<0$.
@@
We also deal with the derivatives of $D_{n}(t)$ and $G_{m}(t)$ with respect to $t$. The following corresponds to [3, (3.32)].
Lemma $3.$
@@
$\lim_{n\to\infty}\max_{t\in[-r_{m,2},r_{m,2}]}\left|D_{n}^{\prime}(t)-G_{m}^{\prime}(t)\right|=0,\ \textup{ $\displaystyle P$-a.s.}$
@@
By
(3.1) $\partial_{t}^{2}D(x,t)=\frac{2(x-t)((x-t)^{2}-3)}{(1+(x-t)^{2})^{3}},$
$|\partial_{t}^{2}D(x,t)|\leq 3$, and hence, the map $t\mapsto\partial_{t}D(x,t)$ is Lipschitz continuous with the Lipschitz constant $3$. Let $Y^{\prime}_{i}\coloneqq\partial_{t}D(X_{i},t)-G_{m}^{\prime}(t)$. Since $|\partial_{t}D(X_{i},t)|\leq 1$ and $|G_{m}^{\prime}(t)|\leq 1$, $\left(Y^{\prime}_{i}\right)_{i}$ are i.i.d., $|Y_{i}^{\prime}|\leq 2$ and $E[Y_{i}^{\prime}]=0$. Therefore, we can show this assertion as in the proof of Lemma 3.1. ∎
@@
$P$-a.s. $\omega$, there exists $N_{5}(\omega)\in\mathbb{N}$ such that for every $n>N_{5}(\omega)$, $\left|C_{Q_{n}(\omega)}\cap[-r_{m,2},r_{m,2}]\right|=1$.
@@
in particular, $D_{n}(t)(\omega)$ is strictly decreasing in $t$ on $[-r_{m,2},r_{m,2}]$.
@@
(3.2) $\mathcal{A}_{n}\coloneqq\mathcal{A}_{n,1}\cap\mathcal{A}_{n,2}\cap\mathcal{A}_{n,3}.$
Let
$\widetilde{\mathcal{A}_{i}}\coloneqq\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n,i},\ \ i=1,2,3.$
By Lemma 3.1, $\displaystyle P\left(\widetilde{\mathcal{A}_{1}}\right)=1$. By Lemma 3.3, $\displaystyle P\left(\widetilde{\mathcal{A}_{2}}\right)=1$. By Propositions 2.7 and 3.4, $\displaystyle P\left(\widetilde{\mathcal{A}_{3}}\right)=1$. Since $\displaystyle\widetilde{\mathcal{A}_{1}}\cap\widetilde{\mathcal{A}_{2}}\cap\widetilde{\mathcal{A}_{3}}=\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}$, $\displaystyle P\left(\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}\right)=1$, and in particular, $\displaystyle\lim_{n\to\infty}P(\mathcal{A}_{n})=1$.
@@
(3.3) $\lim_{n\to\infty}P\left(D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right)=\int_{-\infty}^{y}\varphi_{m}(t)dt,$
where $\varphi_{m}$ is the density function of the distribution $\displaystyle N\left(0,\frac{m+1}{m(2m-1)}\right)$.
@@
By symmetry,
(3.4) $E\left[D(X_{1},0)\right]=c_{m}\int_{\mathbb{R}}\frac{x}{(1+x^{2})^{1+m}}dx=0.$
@@
(3.5) $E\left[D(X_{1},0)^{2}\right]=c_{m}\int_{\mathbb{R}}\frac{x^{2}}{(1+x^{2})^{2+m}}dx=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}=\frac{2m-1}{4m(m+1)},$
@@
(3.6) $\sqrt{n}D_{n}\left(0\right)\Rightarrow N\left(0,\frac{2m-1}{4m(m+1)}\right),\ n\to\infty,$
@@
$\displaystyle E\left[\partial_{t}D(X_{1},0)\right]$ $\displaystyle=c_{m}\int_{\mathbb{R}}\frac{x^{2}-1}{(1+x^{2})^{2+m}}dx$
(3.7) $\displaystyle=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}-\frac{B(1/2,m+3/2)}{B(1/2,m-1/2)}=-\frac{2m-1}{2(m+1)}.$
@@
(3.8) $\lim_{n\to\infty}\frac{y}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0)=-\frac{2m-1}{2(m+1)}y,\ \textup{ $\displaystyle P$-a.s.}$
By (3.1),
$\max_{x,t\in\mathbb{R}}\left|\partial_{t}^{2}D(x,t)\right|=\max_{y\in\mathbb{R}}\frac{2|y||y^{2}-3|}{(1+y^{2})^{3}}\eqqcolon C_{1}<\infty.$
@@
(3.9) $\left|D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right|\leq C_{1}\frac{|y|^{2}}{n}.$
@@
(3.10) $\lim_{n\to\infty}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right)=0,\ \textup{ $\displaystyle P$-a.s.}$
By (3.6), (3.8), (3.10) and Slutsky’s lemma,
@@
Thus we see that (3.3) holds and the proof of Theorem 1.2 is completed.
Subsection 8.2 below provides numerical verifications of Theorem 1.2 by using the Kolmogorov–Smirnov distance. Subsection 8.3 below provides confidence intervals of the parameter $\theta$ by using $\hat{\theta}_{n}$.
@@
(iii) We can apply [22, Theorem 5.23]. We confirm the assumptions of the assertion. Let $\mathfrak{m}(x,\theta)\coloneqq-\log(1+(x-\theta)^{2}),\ x,\theta\in\mathbb{R}$. Then the MLE $\hat{\theta}_{n}$ maximizes the map $\theta\mapsto\sum_{i=1}^{n}\mathfrak{m}(X_{i},\theta)$. We see that $\mathfrak{m}\in C^{\infty}(\mathbb{R}^{2})$ and by (2.3), $|\mathfrak{m}(x,\theta_{1})-\mathfrak{m}(x,\theta_{2})|\leq|\theta_{1}-\theta_{2}|$. It holds that $E\left[\mathfrak{m}(X_{1},\theta)\right]=-F_{m}(\theta)$. Since $F_{m}^{\prime}=-2G_{m}$ and the argument in the proof of Lemma 3.2, $G_{m}\in C^{1}(\mathbb{R})$ and hence $F_{m}\in C^{2}(\mathbb{R})$. Hence $F_{m}$ admits the second-order Taylor expansion at $\theta=0$. Furthermore $F_{m}^{\prime\prime}(0)=-2G_{m}^{\prime}(0)>0$ by Lemma 3.2. Finally we recall Theorem $1.$ Thus the assumptions of [22, Theorem 5.23] are satisfied and Theorem 1.2 follows.
@@
We first give an outline of the proof. We follow the strategy of Barczy and Páles [4, Section 5]. As in the case of Theorem 1.2, we cannot apply the result of [4] directly and need to modify several parts. We evaluate the normalized score along the scale of the law of the iterated logarithm and use the decomposition in (4.1) below into the leading fluctuation term $R_{n}$, the drift term $\gamma S_{n}$, and the remainder $T_{n}^{(\gamma)}$. We apply the Kolmogorov law of the iterated logarithm to $R_{n}$, the strong law of large numbers to $S_{n}$, and finally show that $T_{n}^{(\gamma)}$ is negligible.
Let $\phi(n)\coloneqq\sqrt{2n\log\log n}$. Let
$R_{n}\coloneqq\frac{1}{\phi(n)}\sum_{i=1}^{n}D(X_{i},0),$
$S_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0),$
@@
$T_{n}^{(\gamma)}\coloneqq\frac{1}{\phi(n)}\sum_{i=1}^{n}\left(D\left(X_{i},\gamma\frac{\phi(n)}{n}\right)-D(X_{i},0)-\gamma\frac{\phi(n)}{n}\partial_{t}D(X_{i},0)\right).$
Then
(4.1) $R_{n}+\gamma S_{n}+T_{n}^{(\gamma)}=\frac{1}{\phi(n)}\sum_{i=1}^{n}D\left(X_{i},\gamma\frac{\phi(n)}{n}\right).$
Since $\displaystyle\max_{x,t\in\mathbb{R}}|D(x,t)|\leq 1/2$, by the Kolmogorov law of the iterated logarithm, there exists an event $\Omega_{4}$ such that $P(\Omega_{4})=1$ and for every $\omega\in\Omega_{4}$,
(4.2) $\limsup_{n\to\infty}R_{n}(\omega)=\sqrt{\frac{2m-1}{4m(m+1)}}.$
@@
(4.3) $\lim_{n\to\infty}S_{n}(\omega)=-\frac{2m-1}{2(m+1)}.$
Let $\displaystyle\Omega_{6}\coloneqq\Omega_{4}\cap\Omega_{5}\cap\left(\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}\right)$. Then $P\left(\Omega_{6}\right)=1$.
@@
(4.4) $\lim_{n\to\infty}T_{n}^{(\gamma)}(\omega)=0.$
Let $\sigma_{m}\coloneqq\sqrt{\dfrac{m+1}{m(2m-1)}}$. Let $\omega\in\Omega_{6}$ and $\epsilon>0$. Then there exists $N_{9}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{9}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)<(\sigma_{m}+\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\displaystyle\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}+\epsilon)\frac{\phi(n)}{n}\right)<0$ holds. By (4.1), this is equivalent to
(4.5) $R_{n}(\omega)+(\sigma_{m}+\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}+\epsilon)}(\omega)<0.$
@@
(4.6) $\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\leq\sigma_{m}.$
We can show the lower bound in the same manner. There exists $N_{11}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{11}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)>(\sigma_{m}-\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\displaystyle\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}-\epsilon)\frac{\phi(n)}{n}\right)>0$ holds. By (4.1), this is equivalent to
(4.7) $R_{n}(\omega)+(\sigma_{m}-\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}-\epsilon)}(\omega)>0.$
@@
(4.8) $\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\geq\sigma_{m}.$
@@
This completes the proof of Theorem 1.3.
@@
We first give an outline of the proof. We prove Theorem 1.4 by following the strategy of Bai and Fu [3]. We first recall that $\left|P\left(\hat{\theta}_{n}>\epsilon\right)-P(D_{n}(\epsilon)>0)\right|\leq 2P\left(\mathcal{A}_{n}^{c}\right)$ by the arguments of Section 3.
The first step is to show (5.6) below, which states that the probability of the localization event $P\left(\mathcal{A}_{n}^{c}\right)$ decays exponentially fast as $n\to\infty$. In Section 3, we have seen that for $i=1,2$, $P(\mathcal{A}_{n,i}^{c})$ decays exponentially fast. So it remains to control $P\left(\mathcal{A}_{n,3}^{c}\right)$. This is done by two large deviation estimates for $L_{n}(t)$. First, Lemma 5.3 below controls $\displaystyle\inf_{|t|\geq r}L_{n}(t)$ by using Lemmas 5.1 and 5.2 below. Second, Lemma 5.4 below controls $L_{n}(0)$ by an exponential Chebyshev bound. Combining these bounds, we obtain (5.6).
After this, the problem reduces to estimating $P(D_{n}(\epsilon)>0)$. We center $D(X_{i},\epsilon)$ by its mean $G_{m}(\epsilon)$ and then apply two deviation inequalities in Lemmas 5.6 and 5.7 below with its variance $H_{m}(\epsilon)$. The Taylor expansions of $G_{m}$ and $H_{m}$ around $0$ in Lemma 5.5 below together with (5.10) identify the quadratic rate constant, which matches the upper and lower bounds in (1.1) and (1.2) respectively.
Recall the definition of $F_{m}$ in (2.1). Let
$\widetilde{F}_{m}(t)\coloneqq\exp\left(F_{m}(t)\right).$
@@
Let
(5.2) $\lambda_{m}\coloneqq\frac{1}{2}\left(m-\frac{1}{2}\right)$
and
(5.3) $\delta_{m}(r)\coloneqq\frac{F_{m}(r)-F_{m}(0)}{4},\ r>0.$
These definitions will be used frequently not only in this section but also in the following section.
The following corresponds to [3, (3.15)].
@@
Let $r>0$. Assume that $0<\delta<\min\left\{\lambda_{m},\delta_{m}(r)\right\}$. Then there exists a positive constant $c_{m,3}$ depending only on $m$ such that for every $t$ with $|t|>r$ and every $n\geq 1$,
@@
The following proof is similar to the proof of Bernstein’s inequality ([8, Theorem 2.10]). However the estimates are different, see (5.4) below. The proof below is easier in the sense that there is no need to consider the Fenchel–Legendre transform.
@@
We assume that $t>r$. The proof is the same for the case that $t<-r$. We see that
@@
Assume that $0<\lambda<m-\dfrac{1}{2}$. Then
$E\left[\exp\left(\lambda\left|F_{m}(t)-\log(1+(X_{1}-t)^{2})\right|\right)\right]\leq\exp(\lambda F_{m}(t))E\left[\left(1+(X_{1}-t)^{2}\right)^{\lambda}\right]<\infty.$
@@
$E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]=\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right].$
where we let $\displaystyle\Psi(x,t)\coloneqq\log\frac{\widetilde{F}_{m}(t)}{1+(x-t)^{2}}$. Since $E\left[\Psi(X_{1},t)\right]=0$,
$\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right]=1+\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right].$
@@
(5.4) $E\left[\Psi(X_{1},t)^{k}\right]\leq c_{m,4}^{k}P(X_{1}\leq t/2)+F_{m}(t)^{k}P(X_{1}>t/2).$
@@
we see that
$E\left[\Psi(X_{1},t)^{k}\right]\leq c_{m,4}^{k}+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{k},$
where we let $c_{m,5}\coloneqq c_{m}4^{m}$. Hence,
$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right]\leq\frac{\lambda^{2}}{2}\left(c_{m,4}^{2}\exp(\lambda c_{m,4})+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda F_{m}(t))\right).$
Since $0<\lambda<m-\dfrac{1}{2}$,
@@
Recall (5.2). Then, for every $\lambda\in(0,\lambda_{m})$,
$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right]\leq\lambda^{2}c_{m,6},$
@@
$c_{m,6}\coloneqq\frac{1}{2}\left(c_{m,4}^{2}\exp(\lambda_{m}c_{m,4})+\sup_{t\geq 0}\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda_{m}F_{m}(t))\right)<\infty.$
@@
Therefore, for every $\displaystyle\lambda\in\left(0,\min\left\{\lambda_{m},\delta_{m}(r)\right\}\right)$,
@@
If we let $\lambda\coloneqq\delta/c_{m,6}$, then, $0<\lambda<m-\dfrac{1}{2}$, and,
@@
Let $\displaystyle\lambda_{m}(r)\coloneqq\frac{1}{2}\min\left\{\lambda_{m},\delta_{m}(r)\right\}$.
The following corresponds to [3, (3.21)] [^4] There is a typo in [3, (3.21)]. The supremum in [3, (3.21)] should be the infimum..
@@
Let $r>0$. Assume that $\displaystyle 0<\delta<\lambda_{m}(r)$. Then there exists $N(r,\delta)\in\mathbb{N}$ such that for every $n\geq N(r,\delta)$,
@@
We remark that $r>0$ can be taken arbitrarily small. We first discretize $[r,\infty)$ by the Lipschitz continuity of $L_{n}(t)$ and then apply Lemma 5.2.
@@
(5.5) $P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)\leq\exp\left(-\frac{\delta^{2}}{8c_{m,3}}n\right).$
Since $L_{n}^{\prime}(t)=-2D_{n}(t)$ and $|D_{n}(t)|\leq 1/2$, it holds that
@@
The following corresponds to [3, (3.25)] [^5] There is also a typo in [3, (3.25)]. $n^{2}$” in the right hand side of the inequality in [3, (3.25)] should be $n\delta^{2}$”.. Recall the definition of $\lambda_{m}$ in (5.2).